Weatherly Lumber Company processes wood pulp for manufacturing various paper products. The company employs a process costing system for its manufacturing operations. All direct materials are added at the beginning of the process, and conversion costs are incurred uniformly throughout the process. This is the company’s production schedule for May:

Percent Completed
Tons of Pulp Materials Conversion
Work-in-Process Inventory, May 1 2,000 100 % 50 %
Started during May 11,000
Units to account for 13,000
Units from beginning Work-in-Process, which
were completed and transferred out during May
2,000
Started and completed during May 6,600
Work-in-Process Inventory, May 31 4,400 100 % 50 %
Total units accounted for 13,000

The following cost data are available:

Work-in-Process Inventory, May 1
Direct materials $ 38,010
Conversion 46,568
Costs incurred during May
Direct materials 168,300
Conversion 275,380

Required:

1. Calculate the equivalent units of direct materials and conversion during May. Use the FIFO method.

Equivalent Units
Direct Materials
Conversion

2. Calculate the cost per equivalent unit for both direct materials and conversion during May. Use the FIFO method.

Cost per Equivalent Unit
Direct Materials
Conversion
Answer #1
Percent Completed
Tons of Pulp Materials Conversion
Work-in-Process Inventory, May 1 2,000 100 % 50 %
Started during May 11,000
Units to account for 13,000
Units from beginning Work-in-Process, which
were completed and transferred out during May
2,000
Started and completed during May 6,600
Work-in-Process Inventory, May 31 4,400 100 % 50 %
Total units accounted for 13,000

The following cost data are available:

Work-in-Process Inventory, May 1
Direct materials $ 38,010
Conversion 46,568
Costs incurred during May
Direct materials 168,300
Conversion 275,380

Required:

1. Calculate the equivalent units of direct materials and conversion during May. Use the FIFO method.

Equivalent Units
Direct Materials 11000
Conversion (2000*50%+6600+4400*50%) = 9800

2. Calculate the cost per equivalent unit for both direct materials and conversion during May. Use the FIFO method.

Cost per Equivalent Unit
Direct Materials 168300/11000 = 15.30
Conversion 275380/9800 = 28.10